# Cartesian and spherical multipole expansions in anisotropic media

### v3

@@ II.1 Spherical and Cartesian multipole expansions in isotropic media
The next step is to assume that under appropriate boundary conditions (e.g., radiation conditions for the spherical Hankel function of the first or second kind), there are unique solutions $f_{\ell m}(\mathbf{r})$ and $f_{\alpha}(\mathbf{r})$ to the wave equation with sources, equations (1) and (9). Conversely, the Helmholtz operator applied to the solutions $f_{\ell m}(\mathbf{r})$ and $f_{\alpha}(\mathbf{r})$ yields singularities that are entirely determined by the polynomials $\tilde{Y}(\mathbf{k})$ and $\mathbf{k}^{\alpha}$. Hence, from now on, we identify spherical solutions $f_{\ell m}(\mathbf{r})$ by their corresponding harmonic polynomial $\tilde{Y}_{\ell m}(\mathbf{k})$ and the Cartesian solutions $f_{\alpha}(\mathbf{r})$ by the monomials $\mathbf{k}^{\alpha}$:
$f_{\ell m}(\mathbf{r})=\text{vp}\left[h_{\ell}^{(1)}(k_{0}r)\mathcal{Y}_{\ell m}(\theta,\phi)\right]\iff\tilde{Y}(\mathbf{k})$ (13)
@@ II.1.2 Cartesian to spherical multipole expansions
$\tilde{A}_{|\alpha|-2n,m}=\frac{k_{0}^{\ell+1}}{(-1)^{\ell}4\pi\text{i}}A_{|\alpha|-2n,m}$ (23)
@@ II.3.1 From vector spherical harmonics to the Cartesian multipole expansion
The vector spherical harmonics used by Jackson [25] $\bm{\mathcal{Y}}_{\ell m}^{(q)}(\theta,\phi)$, $q=0,\pm 1$, are obtained as linear combinations of $\bm{\mathcal{Y}}^{\ell}_{jm}(\theta,\phi)$ (see supplementary equations (1-3)). They correspond to the azimuthal ($q=0$), polar ($q=1$), and radial ($q=-1$) components. In the context of Mie scattering, $\bm{\mathcal{Y}}_{\ell m}^{(1)}(\theta,\phi)$ represents electric multipoles, whereas $\bm{\mathcal{Y}}_{\ell m}^{(0)}(\theta,\phi)$ corresponds to magnetic multipoles. In homogeneous media, the radial component is not allowed for electromagnetic fields. Together with Equation (31), we show in the Supplementary Method 1 how to express the azimuthal or polar vector spherical harmonics using the Cartesian multipole expansion of an appropriate current source. We show that a current density given by
@@ II.3.2 From the Cartesian multipole expansion to the vector spherical harmonics
Let us now count the vector spherical harmonics. The azimuthal vector harmonics correspond to a current density given by the vector spherical harmonics $\bm{\mathcal{Y}}_{\ell m}^{\ell}(\theta,\phi)$. In light of Equation (31), there are $12\ell+$ 1 different azimuthal waves of order $\ell>0$ (none for $\ell=0$). By the same argument, there are $12(\ell\pm 1)+$ 1 polar vector harmonics of order $\ell>0$ corresponding to $j=\ell\pm 1$. But the polar harmonics are degenerate, in that both values of $j$ yield the same waves, by virtue of the contraction in Equation (26) (see the Supplementary Method 1.2.3). Thus, in total, there are
@@ III Illustration on an inverse problem
To avoid the “inverse crime” [48], the fitting data (i.e., the prediction through a multipole expansion) is obtained from the analytical Green’s function [49]. The Cartesian multipole expansion basis functions are obtained by differentiating Green’s function up to a fixed order $\ell_{\text{max}}$, i.e.,
@@
Figure 1: Flowchart of the inverse problem for $\varepsilon_{zz}=4$ and $\alpha=(0,0,2)$. (a) The ground truth data consists of all components (pictured: norm of the $x$ component, max-normalized; warm colors indicate high values) of the electric field radiated by the source $\mathbf{J}(\mathbf{r})=\nabla^{\alpha}\delta(\mathbf{r})\mathbf{e}_{x}$ simulated using Meep and sampled on a sphere surrounding the source. The explaining data consists of the analytical Cartesian (b) and spherical (c) multipole expansion basis functions. These basis functions are multiplied by coefficients (d) determined by an optimization algorithm (e), whose goal is to minimize the $L^{2}$-norm of the error (f) between the ground truth and the multipole expansion (g). This error is normalized to the ground truth $L^{2}$-norm. Its complement after convergence represents the proportion of ground truth energy explained by the Cartesian (h) and spherical (i) multipole expansions. The fields obtained after convergence of the minimizer are shown in (j) for the Cartesian and (k) for the spherical multipole expansion.
@@
Solving this ordinary differential equation, we obtain (see Supplementary Method 2)

### v4

@@ II.1 Spherical and Cartesian multipole expansions in isotropic media
The next step is to assume that under appropriate boundary conditions (e.g., radiation conditions for the spherical Hankel function of the first or second kind), there are unique solutions $f_{\ell m}(\mathbf{r})$ and $f_{\alpha}(\mathbf{r})$ to the wave equation with sources, equations (1) and (9). Conversely, the Helmholtz operator applied to the solutions $f_{\ell m}(\mathbf{r})$ and $f_{\alpha}(\mathbf{r})$ yields singularities that are entirely determined by the polynomials $\tilde{Y}_{\ell m}(\mathbf{k})$ and $\mathbf{k}^{\alpha}$. Hence, from now on, we identify spherical solutions $f_{\ell m}(\mathbf{r})$ by their corresponding harmonic polynomial $\tilde{Y}_{\ell m}(\mathbf{k})$ and the Cartesian solutions $f_{\alpha}(\mathbf{r})$ by the monomials $\mathbf{k}^{\alpha}$:
$f_{\ell m}(\mathbf{r})=\text{vp}\left[h_{\ell}^{(1)}(k_{0}r)\mathcal{Y}_{\ell m}(\theta,\phi)\right]$ $\iff\tilde{Y}_{\ell m}(\mathbf{k})$ (13)
@@ II.1.2 Cartesian to spherical multipole expansions
$\tilde{A}_{|\alpha|-2n,m}=\frac{k_{0}^{|\alpha|-2n+1}}{(-1)^{|\alpha|-2n}4\pi\text{i}}A_{|\alpha|-2n,m}$ (23)
@@ II.3.1 From vector spherical harmonics to the Cartesian multipole expansion
The vector spherical harmonics used by Jackson [25] $\bm{\mathcal{Y}}_{\ell m}^{(q)}(\theta,\phi)$, $q=0,\pm 1$, are obtained as linear combinations of $\bm{\mathcal{Y}}^{\ell}_{jm}(\theta,\phi)$ (see supplementary equations (S1-S3)). They correspond to the azimuthal ($q=0$), polar ($q=1$), and radial ($q=-1$) components. In the context of Mie scattering, $\bm{\mathcal{Y}}_{\ell m}^{(1)}(\theta,\phi)$ represents electric multipoles, whereas $\bm{\mathcal{Y}}_{\ell m}^{(0)}(\theta,\phi)$ corresponds to magnetic multipoles. In homogeneous media, the radial component is not allowed for electromagnetic fields. Together with Equation (31), we show in the Supplementary Note 1 how to express the azimuthal or polar vector spherical harmonics using the Cartesian multipole expansion of an appropriate current source. We show that a current density given by
@@ II.3.2 From the Cartesian multipole expansion to the vector spherical harmonics
Let us now count the vector spherical harmonics. The azimuthal vector harmonics correspond to a current density given by the vector spherical harmonics $\bm{\mathcal{Y}}_{\ell m}^{\ell}(\theta,\phi)$. In light of Equation (31), there are $12\ell+$ 1 different azimuthal waves of order $\ell>0$ (none for $\ell=0$). By the same argument, there are $12(\ell\pm 1)+$ 1 polar vector harmonics of order $\ell>0$ corresponding to $j=\ell\pm 1$. But the polar harmonics are degenerate, in that both values of $j$ yield the same waves, by virtue of the contraction in Equation (26) (see the Supplementary Note 1.2.3). Thus, in total, there are
@@ III Illustration on an inverse problem
To avoid the “inverse crime,” [48] the fitting data (i.e., the prediction through a multipole expansion) is obtained from the analytical Green’s function [49]. The Cartesian multipole expansion basis functions are obtained by differentiating Green’s function up to a fixed order $\ell_{\text{max}}$, i.e.,
@@
Figure 1:
Flowchart of the inverse problem for $\varepsilon_{zz}=4$ and $\alpha=(0,0,2)$. (a) The ground truth data consists of all components (pictured: norm of the $x$ component, max-normalized; warm colors indicate high values) of the electric field radiated by the source $\mathbf{J}(\mathbf{r})=\nabla^{\alpha}\delta(\mathbf{r})\mathbf{e}_{x}$ simulated using Meep and sampled on a sphere surrounding the source. The explaining data consists of the analytical Cartesian (b) and spherical (c) multipole expansion basis functions. These basis functions are multiplied by coefficients (d) determined by an optimization algorithm (e), whose goal is to minimize the $L^{2}$-norm of the error (f) between the ground truth and the multipole expansion (g). This error is normalized to the ground truth $L^{2}$-norm. Its complement after convergence represents the proportion of ground truth energy explained by the Cartesian (h) and spherical (i) multipole expansions. The fields obtained after convergence of the minimizer are shown in (j) for the Cartesian and (k) for the spherical multipole expansion.
@@
Solving this ordinary differential equation, we obtain (see Supplementary Note 2)